# CTRL-F: Pairing Convolution with Transformer for Image Classification via Multi-Level Feature Cross-Attention and Representation Learning Fusion

### v1

@@ Abstract
Transformers have captured growing attention in computer vision, thanks to its large capacity and global processing capabilities. However, transformers are data hungry, and their ability to generalize is constrained compared to Convolutional Neural Networks (ConvNets), especially when trained with limited data due to the absence of the built-in spatial inductive biases present in ConvNets. In this paper, we strive to optimally combine the strengths of both convolution and transformers for image classification tasks. Towards this end, we present a novel lightweight hybrid network that pairs Convolution with Transformers via Representation Learning Fusion and Multi-Level Feature Cross-Attention named CTRL-F. Our network comprises a convolution branch and a novel transformer module named multi-level feature cross-attention (MFCA). The MFCA module operates on multi-level feature representations obtained at different convolution stages. It processes small patch tokens and large patch tokens extracted from these multi-level feature representations via two separate transformer branches, where both branches communicate and exchange knowledge through cross-attention mechanism. We fuse the local responses acquired from the convolution path with the global responses acquired from the MFCA module using novel representation fusion techniques dubbed adaptive knowledge fusion (AKF) and collaborative knowledge fusion (CKF). Experiments demonstrate that our CTRL-F variants achieve state-of-the-art performance, whether trained from scratch on large data or even with low-data regime. For Instance, CTRL-F achieves top-1 accuracy of 82.24% and 99.91% when trained from scratch on Oxford-102 Flowers and PlantVillage datasets respectively, surpassing state-of-the-art models which showcase the robustness of our model on image classification tasks. Code at:
@@ 1 Introduction
In this work, we focus how to optimally design a hybrid, resource-efficient network that combines the merits of convolution and transformers. To this end, we introduce a hybrid network named CTRL-F. The network incorporates a CNN pathway, employing inverted residual blocks (MBConv) with squeeze-excitation (SE), alongside a novel Multi-level Feature Cross-Attention Transformer (MFCA) module. Inspired by the success of multi-branch CNN [31, 32] and transformer architectures [17, 33]. Our MFCA module resembles a dual-branch transformer structure, operating on high-level feature maps generated by the CNN pathway. It enriches the network with higher capacity and global processing ability, and hence enhancing its capability of capturing long-range dependencies. The module has two separate branches, each comprising a sequence of transformers encoders, one branch operates on small patch tokens extracted from the latest feature maps produced by the CNN path, while the other branch operates on larger patch tokens extracted from intermediate feature maps. The knowledge gained from each branch are mutually exchanged multiple times through cross-attention layers to complement each other. Finally, we introduce two modules for knowledge fusion: Adaptive Knowledge Fusion (AKF) and Collaborative Knowledge Fusion (CKF). Both modules fuse the local responses acquired from the CNN pathway with the global responses acquired from the MFCA module. The AKF module emphasizes the CNN’s inductive biases in the early training stages and as the training proceeds, it increasingly prioritizes the transformer’s global context. However, The CKF module concatenates the representations generated by the CNN path and the MFCA module efficiently, ensuring the optimization of the entire network through leveraging knowledge from both branches.
@@ 2 Related work
Vision Transformers: Following the remarkable success of transformers in machine translation [5] due to its ability to encode long-term dependencies between tokens, Vision Transformer (ViT) [9] has been introduced to the vision community as the first pure transformer architecture that achieves comparable results with the state-of-the-art CNNs on various benchmark datasets [10, 50, 51]. However, ViTs lack the inductive biases present in CNNs limiting their performance on small datasets. Thus, ViT was initially trained on a large-scale JFT300M dataset [11] to be able to achieve comparable results with CNNs when finetuned on smaller datasets (e.g., ImageNet). Later, DeiT [12] introduces a teacher-student knowledge distillation strategy suitable for transformers through adding an additional distillation token and explores efficient augmentation and regularization techniques for training ViTs. Other approaches [13, 52], effectively extract inductive biases from intermediate layers of a CNN teacher and utilize them as supervisory signals for the transformer student. Further, inspired by the inherent hierarchical structure of convolutional neural networks (CNNs), researchers have explored transformer architectures that incorporate pyramid and hierarchical designs to enhance transformer performance on downstream tasks. PVTv1 [16] introduces a vision transformer with a shrinking pyramid design, while reducing the computational cost and enhancing performance on various dense predictions tasks. Swin Transformer [18] introduces shifted local window attention for capturing long-range dependencies with linear complexity. CrossViT [17] uses dual transformer branches for processing image patches with different resolutions and uses cross-attention for exchanging knowledge between both branches.
@@
Combining CNNs and Transformers: Several recent works unravel how to effectively integrate the complementary strengths of ConvNets and Transformers into a unified hybrid model. Conformer [24] employs dual structure network of CNN and Transformer for local and global features fusion. CvT [53], LeViT [54] and ViTC [22] uses early convolution to replace the regular ViT’s patch embedding for better feature representation. MobileFormer [25] introduces a lightweight network that parallelizes MobileNetV2 [39] and ViT [9] with a two-way bridge connection enabling bidirectional local-global feature fusion. PVTv2 [55] reduces the computational cost over PVTv1 [16] by employing a linear spatial reduction attention. In addition, it uses an overlapping patch embedding and convolutional feed-forward network for enhanced feature representation. MobileViT [26] introduces a light-weight MobileViT block that combines convolution with transformer for modeling local-global representations. MobileViTv2 [27] replaces the multi-headed self-attention (MHA) in MobileViT with separable self-attention for linear complexity. CoAtNet [28] sequentially utilizes MBConv blocks and transformer blocks with relative attention. MOAT [29] effectively merge the MBConv block along with the self-attention in one block rather than using them sequentially.
@@ 3.1 Overview
MBConv Block. is one of the most widely used blocks recognized for its efficiency, lightweight design, and is sometimes termed as ”inverted residual block”. MBConv block initially expands the input feature channels with a factor of 4 through a 1x1 convolution, then a 3x3 depthwise convolution is applied to efficiently grasp the spatial interactions. A Squeeze-Excitation (SE) block [44] is then employed to extract per-channel global information for channel-wise features adaptive recalibration. Finally, a 1x1 convolution is used to project back the features into the original size by reducing the channels with a factor of 4. Furthermore, a residual connection is incorporated for enabling the flow of information between the input and the output. MBConv block is depicted in Figure 1.
Transformer Block. It is used as a core building block in transformer models. It excels at capturing long-range dependencies and relationships between image patches. The transformer encoder consists of a stack of encoder blocks, each with two sub-layers: multiheaded self-attention (MSA) and feed-forward network (FFN). The MSA computes the pairwise similarities between image patches across multiple heads, where each head focuses on capturing different aspects of the relationships between patches. The FFN consists of two fully connected layers separated by a GeLU [56] non-linearity. The first fully connected layer expands the input dimension with a factor , and the second layer shrinks the input dimension back to the original dimension. Residual connection is applied after each block, and layer normalization (LN) is applied before each block. The transformer processing blocks can be expressed as follows:
@@ 3.3 Multi-level Feature Cross-Attention Transformer (MFCA) module
The MFCA module depicted in Figure 3, utilizes a dual-branch transformer structure like CrossViT [17] but the main difference that it operates on high-level feature maps produced from the CNN pathway. The MFCA module take advantage of operating on high level, low-resolution feature maps, serving two main purposes. First, it operates on enhanced feature representations, and benefits from leveraging the inductive biases acquired through convolution. Second, it minimizes the computational overhead by processing low-resolution feature maps, avoiding the high complexity of directly processing the original image pixels, making the network feasible in practice for real-time use.
@@ 3.4 Knowledge fusion
To effectively integrate the local and global information learned from both branches (i.e., convolutional path and MFCA module), we introduce two knowledge fusion techniques named adaptive knowledge fusion and collaborative knowledge fusion. These proposed techniques are designed to guarantee that the knowledge gained from both branches significantly contribute to the final prediction of the model, avoiding any noticeable bias towards a particular branch during the training process.
@@ 3.4.1 Adaptive knowledge fusion
The proposed adaptive knowledge fusion technique is depicted in Figure 5. When using this fusion technique, we let the final linear layer in the CNN branch and MFCA module produce a vector with dimensions corresponding to the number of classes (). The two output vectors representing the CNN prediction () and the transformer prediction () undergo L1-normalization. Subsequently, the normalized CNN prediction is fused with the normalized transformer prediction using an adaptive weighting hyperparameter () to obtain the final prediction of our proposed network. The lambda hyper-parameter is dynamically adapted to prioritize the CNN prediction in the initial epochs and as the training progress, it begins increasingly to pay more attention for the global responses by adapting the lambda gradually to focus more on the transformer prediction. Prioritizing the CNN prediction in the initial training stages also ensures that the MFCA module receives high-quality feature maps, enabling the transformer to effectively extract and utilize global context. Its noteworthy to mention that it was observed that without applying L1-normalization to both output vectors, the model tends to favor the CNN prediction by boosting the magnitude of its vector elements heavily over the transformer prediction leading to diminishing the impact of transformer during training. This challenge arises because the model, in its pursuit of immediate optimization gains, leans heavily towards following the guidance from the CNN and doesn’t account for the long-term contribution of the transformer which will lead to a better global optimization at the late training stages. Hence, normalizing both branches’ predictions is crucial as it forces the model to learn from both branches since both vectors are mapped to comparable range of values (-1 to 1), ensuring that the final prediction can benefit from both branches’ capabilities. The final prediction is given as follows:
where is an adaptive weighting hyperparameter that decreases uniformly over the number of training epochs (i.e., 0.7 → 0.3), and is a scaling hyperparameter that enlarge the fused normalized vector for better convergence.
@@ 4 Experiments
In this section, we perform extensive experiments to demonstrate the efficiency of our proposed novel CTRL-F models compared to some state-of-the-art Convolutional Neural Nets (ConvNets), Vision Transformers (ViTs), and hybrid models across different datasets.
@@ 4.1 Experimental Setup
Datasets. We conducted our experiments on two benchmarks datasets: PlantVillage dataset [58] and Oxford-102 Flowers dataset [51]. The PlantVillage dataset is widely recognized as the largest open-source dataset for classifying various plant leaf diseases. It consists of 54303 healthy and diseased leaf images belonging to 38 categories from 14 distinct plant crop species. For training, 80% of images from each category were used, while the remaining 20% are used for evaluation. The Oxford-102 Flowers dataset comprises 102 unique flower categories, each containing between 40 and 258 images. The training set consists of only 20 images per class, while the remaining images are used for evaluation. This dataset poses a significant challenge due to the limited number of images available for training in each class. The distribution of training and testing sets for each dataset is given in Table 1.
@@
Implementation Details. The proposed CTRL-F models are trained from scratch on both the PlantVillage and Oxford-102 datasets. Our models are trained for 100 epochs on PlantVillage dataset and 200 epochs on Oxford-102 dataset. We utilize an image resolution of 224 224 and employ a batch size of 32 during the training process. The models are trained using PyTorch on NVIDIA P100 GPU using AdamW optimizer [59] with a cosine decay learning rate scheduler. We used an initial learning rate of 5e-4, a minimum learning rate of 1e-5, a weight decay of 0.05 and 5 epochs of linear warm-up. In addition, basic data augmentation (e.g., horizontal flipping and rotation) are used during training.
@@ 4.2 Ablation Studies
Comparison of the proposed knowledge fusion strategies. We compare our small (CTRLF-S) and base (CTRLF-B) variants under the proposed adaptive knowledge fusion (AKF) and collaborative knowledge fusion (CKF) strategies (see section 3.4). Both fusion strategies demonstrate superior performance with the proposed variants when trained from scratch on both datasets. Table 3 presents the top-1 accuracy and the number of parameters for each variant when applying each strategy. For the Oxford-102 dataset, employing the adaptive knowledge fusion strategy with the base variant (CTRLF-B + AKF) achieves top-1 accuracy of 82.24%, surpassing its collaborative knowledge fusion counterpart (CTRLF-B + CKF) with an absolute gain of 0.78%. Whereas, employing the adaptive knowledge fusion strategy with the small variant (CTRLF-S + AKF) achieves top-1 accuracy of 80.79%, surpassing its collaborative knowledge fusion counterpart (CTRLF-S + CKF) with an absolute gain of 1.56%. However, for the PlantVillage dataset, the CTRLF-B + CKF outperforms CTRLF-B + AKF with an absolute gain of 0.02%, while the CTRLF-S + CKF outperforms CTRLF-S + AKF with an absolute gain of 0.06%. Despite the base variant has about 2 number of parameters and FLOPs compared to the small variant, it is observed that the base variants outperform the small variant with 1.45% and 2.23% when using AKF and CKF knowledge fusion techniques respectively, which proves the scalability of our CTRL-F.
Variant | Knowledge | Params | Oxford-102 | PlantVillage |
fusion | top-1 acc | top-1 acc | ||
CTRLF-S | AKF | 9.97M | 80.79% | 99.85% |
CTRLF-S | CKF | 9.99M | 79.23% | 99.91% |
CTRLF-B | AKF | 21.36M | 82.24% | 99.89% |
CTRLF-B | CKF | 21.39M | 81.46% | 99.91% |
Investigating the individual contribution of each branch after training the entire architecture. Since the adaptive knowledge fusion (AKF) strategy combines the convolutional path prediction with the MFCA module prediction using an adaptive weighting hyperparameter ( to obtain the networks’ final prediction, we can easily investigate the impact of each branch prediction on its own and how their ensemble enhances the whole network predictive capabilities. Table 4 reports the top-1 accuracy on both datasets for the small variant when using the CNN prediction only and when using the MFCA module prediction only, and when combining both using the proposed adaptive knowledge fusion strategy. It can be observed that combining both predictions yields favorable performance, ensuring that the model can benefit from effectively fusing local-global representations.
@@
In addition, the selected intermediate feature maps processed by the large transformer branch in the MFCA module has 4 resolution of the final feature maps processed by the small transformer branch. Thus, selecting a patch size of 8 for the large branch and 2 for the small branch enables each transformer encoder to process patches holding the same content but with different level of granularity. This setup is necessary for exchanging the knowledge in cross-attention between patches pointing to the same receptive field but with different representations, ensuring efficient exchange between multi-level feature representations. For instance, allowing the large branch to process patches of size 88, and the small branch to process patches of size 22 surpasses other choices of patch sizes in terms of accuracy when using the AKF and CKF knowledge fusion techniques. Moreover, the AKF knowledge fusion technique with these patch sizes (88 for the large branch and 22 for the small branch) achieves comparable accuracy to using a 22 patch size for both branches but with much fewer FLOPs, as given in Table 5.
@@
CKF | 8 | 2 | 1.43G | 9.99M | 79.23% |
2 | 2 | 3.75G | 9.20M | 79.05% | |
8 | 7 | 1.37G | 11.46M | 78.91% | |
14 | 2 | 1.33G | 12.15M | 79.23% | |
AKF | 8 | 2 | 1.43G | 9.97M | 80.79% |
2 | 2 | 3.75G | 9.18M | 80.74% | |
8 | 7 | 1.37G | 11.44M | 80.28% | |
14 | 2 | 1.33G | 12.12M | 80.64% |
Impact of image resolution on model performance. Since Oxford-102 dataset has variable high-resolution images. We investigate the effect of training the CTRL-F variants when using larger image resolution for Oxford-102 dataset. It is observed from Table 6 that our model demonstrates improved performance by operating at higher resolutions. This enhancement in performance underscores the robustness and adaptability of our model.
Variant | Image size | FLOPs | Accuracy |
CTRLF-S (+CKF) | 2242 | 1.43G | 79.23% |
3842 | 4.19G | 80.29% | |
CTRLF-S (+AKF) | 2242 | 1.43G | 80.79% |
3842 | 4.19G | 83.36% | |
CTRLF-B (+CKF) | 2242 | 3.29G | 81.46% |
3842 | 9.65G | 82.86% | |
CTRLF-B (+AKF) | 2242 | 3.29G | 82.24% |
3842 | 9.65G | 84.06% |

### v2

@@ Abstract
Transformers have captured growing attention in computer vision, thanks to its large capacity and global processing capabilities. However, transformers are data hungry, and their ability to generalize is constrained compared to Convolutional Neural Networks (ConvNets), especially when trained with limited data due to the absence of the built-in spatial inductive biases present in ConvNets. In this paper, we strive to optimally combine the strengths of both convolution and transformers for image classification tasks. Towards this end, we present a novel lightweight hybrid network that pairs Convolution with Transformers via Representation Learning Fusion and Multi-Level Feature Cross-Attention named CTRL-F. Our network comprises a convolution branch and a novel transformer module named multi-level feature cross-attention (MFCA). The MFCA module operates on multi-level feature representations obtained at different convolution stages. It processes small patch tokens and large patch tokens extracted from these multi-level feature representations via two separate transformer branches, where both branches communicate and exchange knowledge through cross-attention mechanism. We fuse the local responses acquired from the convolution path with the global responses acquired from the MFCA module using novel representation fusion techniques dubbed adaptive knowledge fusion (AKF) and collaborative knowledge fusion (CKF). Experiments demonstrate that our CTRL-F variants achieve state-of-the-art performance, whether trained from scratch on large data or even with low-data regime. For Instance, CTRL-F achieves top-1 accuracy of 84.01%, 61.68% and 99.91% when trained from scratch on Oxford-102 Flowers, CUB-200 and PlantVillage datasets respectively, surpassing state-of-the-art models which showcase the robustness of our model on image classification tasks.
@@ 1 Introduction
In this work, we focus how to optimally design a hybrid, resource-efficient network that combines the merits of convolution and transformers. To this end, we introduce a hybrid network named CTRL-F. The network incorporates a CNN pathway, employing inverted residual blocks (MBConv) with squeeze-excitation (SE), alongside a novel Multi-level Feature Cross-Attention Transformer (MFCA) module. Inspired by the success of multi-branch CNN [31, 32] and transformer architectures [17, 33]. Our MFCA module resembles a dual-branch transformer structure, operating on high-level feature maps generated by the CNN pathway. It enriches the network with higher capacity and global processing ability, and hence enhancing its capability of capturing long-range dependencies. The module has two separate branches, each comprising a sequence of transformers encoders, one branch operates on small patch tokens extracted from the latest feature maps produced by the CNN path, while the other branch operates on larger patch tokens extracted from intermediate feature maps. The knowledge gained from each branch is mutually exchanged multiple times through cross-attention layers to complement each other. Finally, we introduce two modules for knowledge fusion: Adaptive Knowledge Fusion (AKF) and Collaborative Knowledge Fusion (CKF). Both modules fuse the local responses acquired from the CNN pathway with the global responses acquired from the MFCA module. The AKF module emphasizes the CNN’s inductive biases in the early training stages and as the training proceeds, it increasingly prioritizes the transformer’s global context. However, The CKF module concatenates the representations generated by the CNN path and the MFCA module efficiently, ensuring the optimization of the entire network through leveraging knowledge from both branches.
@@ 2 Related work
Vision Transformers: Following the remarkable success of transformers in machine translation [5] due to its ability to encode long-term dependencies between tokens, Vision Transformer (ViT) [9] has been introduced to the vision community as the first pure transformer architecture that achieves comparable results with the state-of-the-art CNNs on various benchmark datasets [10, 50, 51]. However, ViTs lack the inductive biases present in CNNs limiting their performance on small datasets. Thus, ViT was initially trained on a large-scale JFT300M dataset [11] to be able to achieve comparable results with CNNs when finetuned on smaller datasets (e.g., ImageNet). Later, DeiT [12] introduces a teacher-student knowledge distillation strategy suitable for transformers through adding an additional distillation token and explores efficient augmentation and regularization techniques for training ViTs. Other approaches [13, 52], effectively extract inductive biases from intermediate layers of a CNN teacher and utilize them as supervisory signals for the transformer student. Further, inspired by the inherent hierarchical structure of convolutional neural networks (CNNs), researchers have explored transformer architectures that incorporate pyramid and hierarchical designs to enhance transformer performance on downstream tasks. PVTv1 [16] introduces a vision transformer with a shrinking pyramid design, while reducing the computational cost and enhancing performance on various dense predictions tasks. Swin Transformer [18] introduces shifted local window attention for capturing long-range dependencies with linear complexity. CrossViT [17] uses dual transformer branches for processing raw image patches with different resolutions and uses cross-attention for exchanging knowledge between both branches.
@@
Combining CNNs and Transformers: Several recent works unravel how to effectively integrate the complementary strengths of ConvNets and Transformers into a unified hybrid model. Conformer [24] employs dual structure network of CNN and Transformer for local and global features fusion. CvT [53], LeViT [54] and ViTC [22] uses early convolution to replace the regular ViT’s patch embedding for better feature representation. MobileFormer [25] introduces a lightweight network that parallelizes MobileNetV2 [39] and ViT [9] with a two-way bridge connection enabling bidirectional local-global feature fusion. PVTv2 [55] reduces the computational cost over PVTv1 [16] by employing a linear spatial reduction attention. In addition, it uses an overlapping patch embedding and convolutional feed-forward network for enhanced feature representation. MobileViT [26] introduces a light-weight MobileViT block that combines convolution with transformer for modeling local-global representations. MobileViTv2 [27] replaces the multi-headed self-attention (MHA) in MobileViT with separable self-attention for linear complexity. CoAtNet [28] sequentially utilizes MBConv blocks and transformer blocks with relative attention. MOAT [29] effectively merges the MBConv block along with the self-attention in one block rather than using them sequentially.
FasterViT [56] utilizes traditional convolutional blocks in the initial stages and introduces a lightweight module termed Hierarchical Attention block in the later stages.
Building on existing hybrid approaches, our work introduces CTRL-F, which combines lightweight CNN blocks with a Multi-level Feature Cross Attention Transformer module. This design fuses local responses from the CNN and global responses from the transformer using dual fusion approaches, resulting in robust generalization even with limited data. Since CNNs include inductive biases which make it the preferred choice on limited data and Vision transformers include powerful processing capabilities, we combine benefits of both approaches. Moreover, by applying cross attention to high-level CNN features across multiple scales, CTRL-F captures richer inter-scale semantic relationships rather than relying solely on self-attention within individual feature scales.
@@ 3.1 Overview
MBConv Block. is one of the most widely used blocks recognized for its efficiency, lightweight design, and is sometimes termed as ”inverted residual block” [39]. MBConv block initially expands the input feature channels with a factor of 4 through a 1x1 convolution, then a 3x3 depthwise convolution is applied to efficiently grasp the spatial interactions. A Squeeze-Excitation (SE) block [44] is then employed to extract per-channel global information for channel-wise features adaptive recalibration. Finally, a 1x1 convolution is used to project back the features into the original size by reducing the channels with a factor of 4. Batch normalization is applied after each convolution layer and GeLU activation is used. Furthermore, a residual connection is incorporated for enabling the flow of information between the input and the output. MBConv block is depicted in Figure 1.
Transformer Block. It is used as a core building block in transformer models [5]. It excels at capturing long-range dependencies and relationships between image patches. The transformer encoder consists of a stack of encoder blocks, each with two sub-layers: multiheaded self-attention (MSA) and feed-forward network (FFN). The MSA computes the pairwise similarities between image patches across multiple heads, where each head focuses on capturing different aspects of the relationships between patches. The FFN consists of two fully connected layers separated by a GeLU [57] non-linearity. The first fully connected layer expands the input dimension with a factor , and the second layer shrinks the input dimension back to the original dimension. Residual connection is applied after each block, and layer normalization (LN) is applied before each block. The transformer processing blocks can be expressed as follows:
@@ 3.3 Multi-level Feature Cross-Attention Transformer (MFCA) module
The MFCA module depicted in Figure 3, utilizes a dual-branch transformer structure like CrossViT [17] but the main difference that it operates on high-level feature maps produced from the CNN pathway. The MFCA module take advantage of operating on high level, low-resolution feature maps, serving two main purposes. First, it operates on enhanced feature representations, and benefits from leveraging the inductive biases acquired through convolution. Second, it minimizes the computational overhead by processing low-resolution feature maps with better semantic information, avoiding the high complexity of directly processing the original image pixels, making the network feasible in practice for real-time use.
@@ 3.4 Knowledge fusion
To effectively integrate the local and global information learned from both branches (i.e., convolutional path and MFCA module), we introduce two knowledge fusion techniques named adaptive knowledge fusion and collaborative knowledge fusion. These proposed techniques are designed to guarantee that the knowledge gained from both branches significantly contributes to the final prediction of the model, avoiding any noticeable bias towards a particular branch during the training process.
@@ 3.4.1 Adaptive knowledge fusion
The proposed adaptive knowledge fusion technique is depicted in Figure 5. When using this fusion technique, we let the final linear layer in the CNN branch and MFCA module produce a vector with dimensions corresponding to the number of classes (). The two output vectors representing the CNN prediction () and the transformer prediction () undergo L1-normalization. Subsequently, the normalized CNN prediction is fused with the normalized transformer prediction using an adaptive weighting hyperparameter () to obtain the final prediction of our proposed network. The lambda hyper-parameter is dynamically adapted to prioritize the CNN prediction in the initial epochs and as the training progress, it begins increasingly to pay more attention for the global responses by adapting the lambda gradually to focus more on the transformer prediction. Prioritizing the CNN prediction in the initial training stages also ensures that the MFCA module receives high-quality feature maps, enabling the transformer to effectively extract and utilize global context. It’s noteworthy to mention that it was observed that without applying L1-normalization to both output vectors, the model tends to favor the CNN prediction by boosting the magnitude of its vector elements heavily over the transformer prediction leading to diminishing the impact of transformer during training. This challenge arises because the model, in its pursuit of immediate optimization gains, leans heavily towards following the guidance from the CNN and doesn’t account for the long-term contribution of the transformer which will lead to better global optimization at the late training stages. Hence, normalizing both branches’ predictions is crucial as it forces the model to learn from both branches since both vectors are mapped to comparable range of values (-1 to 1), ensuring that the final prediction can benefit from both branches’ capabilities. The final prediction is given as follows:
where is an adaptive weighting hyperparameter that decreases uniformly over the number of training epochs (i.e., 0.7 → 0.3), and is a scaling hyperparameter that enlarges the fused normalized vector for better convergence.
@@ 4 Experiments
In this section, we perform extensive experiments to demonstrate the efficiency of our proposed novel CTRL-F models compared to some state-of-the-art Convolutional Neural Nets (ConvNets), Vision Transformers (ViTs), and hybrid models discussed in the related work section across multiple datasets.
@@ 4.1 Experimental Setup
Datasets. Datasets. We conducted our experiments on three benchmarks datasets: PlantVillage [59], Oxford-102 Flowers [51] and CUB-200 [60] datasets. The PlantVillage dataset is widely recognized as the largest open-source dataset for classifying various plant leaf diseases. It consists of 54303 healthy and diseased leaf images belonging to 38 categories from 14 distinct plant crop species. For training, 80% of images from each category were used, while the remaining 20% are used for evaluation. The Oxford-102 Flowers dataset comprises 102 unique flower categories, each containing between 40 and 258 images. The training set consists of only 20 images per class, while the remaining images are used for evaluation. The CUB-200 dataset is a widely used benchmark for fine-grained visual categorization. It consists of 200 unique bird species, with each class containing approximately 60 images, split into for training and for testing. Both Oxford-102 and CUB-200 datasets pose a significant challenge due to the limited number of images available for training in each class. The distribution of training and testing sets for each dataset is given in Table 1.
@@
| CUB-200 [60] | 5,994 | 5,794 | 200 |
Implementation Details. The proposed CTRL-F models are trained from scratch on PlantVillage, Oxford-102 and CUB-200 datasets. Our models are trained for 100 epochs on PlantVillage dataset and 200 epochs on Oxford-102 and CUB-200 datasets. We utilize an image resolution of 224 224 and employ a batch size of 32 during the training process. The models are trained using PyTorch on NVIDIA P100 GPU using AdamW optimizer [61] with a cosine decay learning rate scheduler. We used an initial learning rate of 5e-4, a minimum learning rate of 1e-5, a weight decay of 0.05 and 5 epochs of linear warm-up. In addition, basic data augmentation (e.g., horizontal flipping and rotation) are used during training. It is noteworthy to mention that the learning rate parameters should be carefully selected to ensure balanced training of both branches (i.e., CNN head and transformer head), preventing one from dominating while the other vanishes.
@@ 4.2 Ablation Studies
Comparison of the proposed knowledge fusion strategies. We compare our small (CTRLF-S) and base (CTRLF-B) variants under the proposed adaptive knowledge fusion (AKF) and collaborative knowledge fusion (CKF) strategies (see section 3.4). Both fusion strategies demonstrate superior performance with the proposed variants when trained from scratch on both datasets. Table 3 presents the top-1 accuracy and the number of parameters for each variant when applying each strategy. For the Oxford-102 dataset, employing the adaptive knowledge fusion strategy with the base variant (CTRLF-B + AKF) achieves top-1 accuracy of 84.01%, surpassing its collaborative knowledge fusion counterpart (CTRLF-B + CKF) with an absolute gain of 1.01%. Whereas, employing the adaptive knowledge fusion strategy with the small variant (CTRLF-S + AKF) achieves top-1 accuracy of 82.42%, surpassing its collaborative knowledge fusion counterpart (CTRLF-S + CKF) with an absolute gain of 0.54%. For the CUB-200 dataset, employing the CTRLF-B + CKF outperforms CTRLF-B + AKF with an absolute gain of 0.91%, while the CTRLF-S + CKF outperforms CTRLF-S + AKF with an absolute gain of 1.08%. However, for the PlantVillage dataset, the CTRLF-B + CKF outperforms CTRLF-B + AKF with an absolute gain of 0.02%, while the CTRLF-S + CKF outperforms CTRLF-S + AKF with an absolute gain of 0.06%. Despite the base variant has about number of parameters and FLOPs compared to the small variant, it is observed that the base variants outperform the small variant when using AKF and CKF knowledge fusion techniques on the three datasets, which proves the scalability of our CTRL-F.
| Variant | Knowledge | Params | Oxford-102 | CUB-200 | PlantVillage |
| fusion | top-1 acc | top-1 acc | top-1 acc | ||
| CTRLF-S | AKF | 9.97M | 82.42% | 56.27% | 99.85% |
| CTRLF-S | CKF | 9.99M | 81.88% | 57.35% | 99.91% |
| CTRLF-B | AKF | 21.36M | 84.01% | 60.77% | 99.89% |
| CTRLF-B | CKF | 21.39M | 83.00% | 61.68% | 99.91% |
Investigating the individual contribution of each branch after training the entire architecture. Since the adaptive knowledge fusion (AKF) strategy combines the convolutional path prediction with the MFCA module prediction using an adaptive weighting hyperparameter ( to obtain the networks’ final prediction, we can easily investigate the impact of each branch prediction on its own and how their ensemble enhances the whole network predictive capabilities. Table 4 reports the top-1 accuracy for AKF strategy on the experimented datasets for the small variant when using the CNN head prediction only and when using the MFCA module head prediction only, and when combining both using the proposed adaptive knowledge fusion strategy. It can be observed that combining both predictions yields favorable performance, ensuring that the model can benefit from effectively fusing local-global representations.
@@
In addition, the selected intermediate feature maps processed by the large transformer branch in the MFCA module has 4 resolution of the final feature maps processed by the small transformer branch. Thus, selecting a patch size of 8 for the large branch and 2 for the small branch enables each transformer encoder to process patches holding the same content but with different level of granularity. This setup is necessary for exchanging the knowledge in cross-attention between patches pointing to the same receptive field but with different representations, ensuring efficient exchange between multi-level feature representations. For instance, allowing the large branch to process patches of size 88, and the small branch to process patches of size 22 surpasses other choices of patch sizes in terms of accuracy when using the AKF and CKF knowledge fusion techniques.
@@
| CKF | 8 | 2 | 1.43G | 9.99M | 81.88% |
| 2 | 2 | 3.75G | 9.20M | 80.61% | |
| 8 | 7 | 1.37G | 11.46M | 80.48% | |
| 14 | 2 | 1.33G | 12.15M | 81.47% | |
| AKF | 8 | 2 | 1.43G | 9.97M | 82.42% |
| 2 | 2 | 3.75G | 9.18M | 82.35% | |
| 8 | 7 | 1.37G | 11.44M | 81.80% | |
| 14 | 2 | 1.33G | 12.12M | 82.42% | |
Impact of image resolution on model performance. Since Oxford-102 and CUB-200 datasets have variable high-resolution images. We investigate the effect of training the CTRL-F variants when using larger image resolution for both dataset. It is observed from Table 6 that our model demonstrates improved performance by operating at higher resolutions. This enhancement in performance underscores the robustness and adaptability of our model.
| Variant | Image size | FLOPs | Oxford-102 | CUB-200 |
| top-1 acc | top-1 acc | |||
| CTRLF-S (+CKF) | 2242 | 1.43G | 81.88% | 57.35% |
| 3842 | 4.19G | 82.57% | 63.19% | |
| CTRLF-S (+AKF) | 2242 | 1.43G | 82.42% | 56.27% |
| 3842 | 4.19G | 83.77% | 61.68% | |
| CTRLF-B (+CKF) | 2242 | 3.29G | 83.00% | 61.68% |
| 3842 | 9.65G | 85.56% | 65.48% | |
| CTRLF-B (+AKF) | 2242 | 3.29G | 84.01% | 60.77% |
| 3842 | 9.65G | 85.88% | 65.43% |